# On the Limitations of Fine-tuned Judge Models for LLM Evaluation

### v2

@@ Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have employed proprietary close-source models, especially GPT-4, as the evaluator. Alternatively, other works have fine-tuned judge models based on open-source LLMs as the evaluator. While the fine-tuned judge models are claimed to achieve comparable evaluation capability with GPT-4, in this study, we conduct an empirical study of judge models. Our findings indicate that although the fine-tuned judge models achieve high performance on in-domain test sets, even surpassing GPT-4, they underperform GPT-4 across several dimensions, including generalizability, fairness, aspect-specific evaluation, and scalability. We also reveal that the fine-tuned judge model inherently operates as a task-specific classifier, consequently imposing the limitations. Finally, we propose an effective indicator to measure the reliability of fine-tuned judges, with the aim of maximizing their utility in LLM evaluation [^1] Codes are openly available at https://github.com/HuihuiChyan/UnlimitedJudge.
@@ 1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn considerate attention of research community Liang et al. (2022); Chang et al. (2023). As the capabilities of LLMs continue to develop across various tasks, it is essential to evaluate them from comprehensive perspectives Qin et al. (2023). However, existing benchmarks, such as MMLU Hendrycks et al. (2021) and BIG-bench bench authors (2023), cannot fully showcase the generative ability of LLMs.
@@
Some research has proposed LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namley utilizing proprietary LLMs, especially GPT-4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, LLMs can leverage their instruction-following ability to provide reliable evaluation, achieving a high agreement rate with human evaluators.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed Zhu et al. (2023b); Wang et al. (2024), relying on open-source foundation models and data constructed from either GPT-4 or human annotation, as shown in Figure 1. These models are validated on their respective meta-evaluation benchmarks, where the finetuned models exhibit performance on par with GPT-3.5 and GPT-4, leading to the affirmation of their evaluation capability.
In this paper, we conduct an empirical study of the evaluation capabilities of different judge models. Our study encompasses a comprehensive comparison across various benchmarks and dimensions. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations compared with close-sourced models in the following aspects:
@@
We think these limitations primarily stem from the fine-tuning process, where the foundation model is transformed into a task-specific classifier overfitted to the fine-tuning data.
Finally, we propose an indicator to quantify the reliability of the fine-tuned judge when applied for LLM evaluation. Our indicator is based on confidence estimation from softmax probability distribution, and we apply calibration to better model the confidence brought by the task-specific fine-tuning process. We validate our indicator for the judge model when applied to different meta-evaluation benchmarks, and managed to select the samples lying inside the scope of fine-tuning, on which the performance of the judge is more reliable.
@@
  • We propose an indicator to quantify the reliability of the fine-tuned judge, which can help to select reliable evaluations;
@@ 2 Related Work
As Large Language Models (LLMs) continue to excel across various tasks and domains, it is essential to design an efficient and effective evaluation method. However, traditional evaluation metrics for generative models, such as BLEU Papineni et al. (2002) and ROUGE Lin (2004), are not expressive enough to recognize the good responses. On the other hand, after the emergence of BERT Kenton and Toutanova (2019), there are also several benchmarks designed to evaluation the capabilities of language models, such as MMLU Hendrycks et al. (2021) and BIG-bench bench authors (2023). But they are mostly formalized as multi-choice selection, which cannot not fully showcase the models’ generation capability.
Recent research has introduced the concept of LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. For example, Li et al. (2023c) constructed a test set containing 805 questions and used the win rate compared with text-davinci-003 as the evaluation result, which is determined by GPT-4. Zheng et al. (2023) developed 80 multi-round test questions covering eight common areas, and then automatically scored the model’s answers using GPT-4. The GPT-4-based evaluator is proven to be high accurate compared with professional human evaluators, and presenting even better consistency and stability compared with human.
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. Therefore, follow-up works suggest fine-tuning language models specialized in evaluations. For instance, PandaLM Wang et al. (2024) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B Touvron et al. (2023) as a judge model. JudgeLM Zhu et al. (2023b) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna Chiang et al. (2023) as a scalable judge model. Auto-J Li et al. (2023a) constructs judgement data upon multiple scenarios to train a generative judge model, which can provide both judgement and critic. Prometheus Kim et al. (2023) defines thousands of evaluation criteria and construct a feedback dataset based on GPT-4, and finetunes a fine-grained judge model.
Despite relying on models with 7B or 13B parameters, these fine-tuned judge models all achieve comparable accuracy with GPT-3.5 or GPT-4 on their respective testsets. However, the evaluation is mostly conducted on the in-domain testsets constructed similarly with the training data. A thorough examination to the evaluation capability of the fine-tuned judges is of urgent demand.
@@
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our empirical study. All the four models are open-source, with their training and test data also publicly released.
@@ 3 How Far can Fine-tuned Judges Go?
In this section, we make a comprehensive empirical study based on four representative fine-tuned judge models as listed in Table 1. Section 3.1 offers a brief introduction of the construction of fine-tuned judges, and the subsequent sections explain their limitations one by one.
@@ 3.1 Preliminary: Fine-tuned LLM Judge
Model JudgeLM-test PandaLM-test Auto-J-test Average
accuracy F1 accuracy F1 agreement
JudgeLM-7B 78.98 68.62 68.17 64.87 46.6 64.58
PandaLM-7B 66.44 56.01 68.97 60.95 40.0 58.47
Auto-J-13B 77.19 60.42 72.27 64.27 54.6 68.02
Prometheus-13B 54.24 50.04 45.25 43.58 47.8 49.10
w/o trans 24.58 23.39 29.03 27.92 16.2 23.26
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 59.88
GPT-4-1106 85.28 76.87 74.07 68.09 56.3 71.88
Table 2: Results of evaluators on pairwise selection. Notice Prometheus can be transformed for pairwise selection by grading two answers twice and compare the scores, therefore we release both results with and without transformation.
3.2 Constrained by Evaluation Scheme
Model Prometheus-test-ind Prometheus-test-ood Average
pearson kendalltau spearman pearson kendalltau spearman
Prometheus-13B 0.864 0.788 0.863 0.869 0.789 0.869 0.867
JudgeLM-7B 0.649 0.647 0.739 0.610 0.602 0.690 0.630
w/o trans 0.398 0.371 0.416 0.384 0.371 0.419 0.391
PandaLM-7B 0.417 0.368 0.423 0.386 0.333 0.383 0.402
Auto-J-13B 0.614 0.526 0.608 0.591 0.504 0.580 0.603
GPT-3.5-0613 0.636 0.536 0.617 0.563 0.453 0.521 0.600
GPT-4-1106 0.742 0.659 0.747 0.743 0.660 0.747 0.743
Table 3: Results of evaluators on pointwise grading. Notice JudgeLM can be transformed for pointwise grading by adding the reference as the first answer, therefore we release both results with and without transformation.
Model MTBench
@@ 3.2 Constrained by Evaluation Scheme
One of the most appealing attribute of LLMs is their generalization ability, enabling them to execute various task defined by various instructions Zhu et al. (2023a). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, etc. Since different judge models are fine-tuned on different schemes, we would like to verify their evaluation capability under the scheme defined by others. Specifically, we apply their publicly released checkpoints, and cross-validate the judge models on each other’s testsets [^2] We make minimal modification to the predefined prompts to adapt the model to different schemes. For detailed prompts please refer to Appendix A.2..
As shown in Table 2 and 3, all four models perform the best on their own training schemes, respectively, with results comparable with GPT-4. However, if we employ a model on an evaluation scheme where it is not trained on, the evaluation performance would drop by a large margin. For example, using a pairwise model (such as PandaLM or JudgeLM) for pointwise grading (such as Prometheus testset), or using a pointwise model (such as Prometheus) for pairwise selection (such as PandaLM or JudgeLM testsets), would all lead to catastrophic performance degradation. On the contrary, close-sourced models such as GPT-3.5 or GPT-4 consistently exhibits superior performance across various evaluation schemes.
We also validate the judge models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset\@footnotemark. As shown in Table 4, while the four models are all trained for single-turn evaluation, they underperforms GPT-4 on MT-bench by a large margin. This demonstrates that the finetuned judge models are overfitted to their respective evaluation schemes and lost their generalibility.
3.3 Biased Towards Superficial Quality
Recently, there has been a lot research on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023b); Saito et al. (2023). To address this issue, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises one natural testset (Natural) and four adversarial testsets (Neighbor, Manual, GPTOut, GPTInst), and the adversarial testsets consist of paired outputs with a correct answer and an incorrect answer with better superficial quality (e.g., more fluent, more verbose, etc.). We evaluate the judge models on LLMBar [^3] The detailed prompts are present in Appendix A.2. as shown in Table 5.
@@ 3.3 Biased Towards Superficial Quality
As can be seen, the fine-tuned judge models achieves a poor results on adversarial testsets, even much worse than random-guess. This notifies that they are severely biased to superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting in the preference to the incorrect answers. On the other hand, GPT-4 does not over-rely on the superficial features and achieves decent accuracy on all the testsets. This notifies that the superior performance of fine-tuned judges on the in-domain testsets may rely on spurious statistical features Niven and Kao (2019), instead of really differentiating good and bad responses.
@@
GPT-3.5-0613 57.50 57.10 62.60 60.27 72.10 72.08 95.10 80.80 98.75 97.54
GPT-4-1106 72.50 72.50 72.00 71.44 84.50 84.78 94.50 82.78 100 100
Table 6: Results of evaluators on aspect-specific evaluation. w/o adapt denotes using the original prompt without adaptation to the specific aspect. As HaluEval and ToxicChat are both binary classification, we apply Auto-J and Prometheus with pointwise grading and conduct grid-search to determine the classification threshold. On the other hand, as SALAD-Bench is pairwise classification, we apply pairwise selection models, namely JudgeLM, PandaLM and Auto-J to select a better response.
We also fine-tune a DeBERTa-based judge model in a classification style on LLMBar (please refer to Section 3.6 for details). It deserves noticing that the DeBERTa-based evaluator also outperforms the LLM-based evaluator by a large margin in terms of fairness. This inspires us that the bias of LLM-based evaluator may come from the casual language modeling process. While the model is trained to generate fluent and verbose responses, it also tends to prefer fluent and verbose response when employed for evaluation, even if it is not aligned with the instruction.
3.4 Incapable of Aspect-specific Evaluation
LLM evaluation covers various aspects such as usefulness, safety, and factuality, and sometimes we are particularly interested in a specific aspect. While previous work primarily assess the evaluation capability of the judge models from a general perspective, we would like to assess them on fine-grained aspects. We select the following three datasets:
@@ 3.4 Incapable of Aspect-specific Evaluation
    HaluEval Li et al. (2023b): This dataset focuses on factuality evaluation. It contains generated and human-annotated hallucinated samples, which lie in three domains: Question-Answering, Summary and Dialogue. Given an instruction-response pair, the evaluator should decide whether the response is hallucinated.
@@
    ToxicChat Lin et al. (2023): This dataset focuses on toxicity evaluation. It contains toxic and non-toxic conversations based on real-world user-AI interactions. Given an instruction-response pair, the evaluator should decide whether the response is toxic.
@@
    SALAD-Bench Li et al. (2024): This dataset focuses on safety evaluation. It contains instructions and responses spanning among different domains and tasks. Given an instruction and a pair of responses, the evaluator should decide which response is safer.
@@
3.5 No Benefits from Prompt Engineering
One of the most appealing features of LLM is it can benefit from delicate prompt engineering. Various strategies have been proposed to improve the LLM’s capability on various tasks, including text evaluation. In this section, we select two representative strategies, namely In-context Learning (ICL) Dong et al. (2023) and Chain-of-Thought Prompting (CoT) Wei et al. (2022), to further improve the evaluation capability of the judge models:

### v3

# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Model is not a General Substitute for GPT-4

@@ Abstract
Recently, there has been a growing trend of utilizing Large Language Model (LLM) to evaluate the quality of other LLMs. Many studies have employed proprietary close-sourced models, especially GPT-4, as the evaluator. Alternatively, other works have fine-tuned judge models based on open-source LLMs as the evaluator. While the fine-tuned judge models are claimed to achieve comparable evaluation capability with GPT-4, in this work, we conduct an empirical study of judge models. Our findings indicate that although the fine-tuned judge models achieve high performance on in-domain test sets, even surpassing GPT-4, they underperform GPT-4 across several dimensions, including generalizability, fairness, aspect-specific evaluation, and scalability. We also reveal that the fine-tuned judge model inherently operates as a task-specific classifier, consequently imposing the limitations. Finally, we introduce a integrated method, leveraging GPT-4 to compensate for the limitations and improve the fine-tuned judges. Experiment results show our method achieves accuracy on par with GPT-4 with only 50% of the API expense. [^1] Codes are openly available at https://github.com/HuihuiChyan/UnlimitedJudge
@@ 1 Introduction
Recently, the evaluation for Large-scale Language Models (LLMs) has drawn significant attention of research community Liang et al. (2022); Chang et al. (2023). As the capabilities of LLMs continue to develop across various tasks, it is essential to evaluate them in a scalable and comprehensive manner Qin et al. (2023). Some research has proposed LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT-4 Achiam et al. (2023), to evaluate the LLM’s response. By defining evaluation schemes in the prompt template, proprietary LLMs can provide an accurate evaluation with a high agreement rate with human evaluators.
@@
However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility. To address these issues, several fine-tuned judge models are proposed Zhu et al. (2023b); Wang et al. (2024); Ke et al. (2024), relying on open-source foundation models and data constructed from either GPT-4 or human annotation, as shown in Figure 1. These models are validated on their respective meta-evaluation benchmarks, where the finetuned models exhibit performance on par with GPT-3.5 and GPT-4, leading to the affirmation of their evaluation capability.
With the prevalence of fine-tuned judge models, in this paper, we conduct an empirical study of their evaluation capabilities. Our study encompasses a comprehensive comparison across various benchmarks and dimensions. Experiment results indicate that while the fine-tuned judge models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations compared with close-sourced models in the following aspects:
@@
We argue that these limitations primarily stem from the fine-tuning process, where the foundation model is transformed into a task-specific classifier overfitted to the fine-tuning data.
In light of the empirical findings, we propose an integrated evaluation method, CascadedEval, to compensate for the limitations with the help of a proprietary model. CascadedEval is based on the confidence indicator derived from softmax probability distribution to assess the confidence of the fine-tuned judge. After that, samples with lower confidence are re-evaluated by a proprietary model, such as GPT-4. Experiment results show our method achieves accuracy on par with GPT-4 on both in-domain and out-of-domain testsets, with only 50% of the API expense.
@@
  • We propose an integrated evaluation method by leveraging GPT-4 to compensate for the limitations of fine-tuned judges;
@@ 2 Related Work
2.1 LLM-as-a-Judge
Benchmark evaluation has long been the default standard in the study of traditional pre-trained models Wang et al. (2018, 2019). Therefore, after the emergence of LLMs, numerous efforts have been made to construct benchmark testsets that cover a variety of tasks for evaluation Srivastava et al. (2023); Zhong et al. (2024). However, questions in these benchmarks are mostly formalized as multi-choice selection or completion, which can not fully showcase the models’ generation capability.
Consequently, recent research has introduced the concept of LLM-as-a-Judge Li et al. (2023c); Zheng et al. (2023), namely utilizing proprietary LLMs, especially GPT4 Achiam et al. (2023), to evaluate the LLM’s response. For example, Li et al. (2023c) constructed a test set containing 805 questions and used the win rate compared with text-davinci-003 as the evaluation result, which is determined by GPT-4. Zheng et al. (2023) developed 80 multi-round test questions covering eight common areas, and then automatically scored the model’s answers using GPT-4. The GPT-4-based evaluator is proven to be high accurate compared with professional human evaluators and presents even better consistency and stability compared with human. However, relying on external API for evaluation may introduce consideration about privacy leakage, and the opacity of API models also challenges the evaluation reproducibility.
@@ 2.1 LLM-as-a-Judge
Table 1: Detailed statistics of the four fine-tuned judge models, which is the foundation of our empirical study. All four models are open-source, with their training and test data also publicly released.
2.2 Finetuned LLM as a Judge
In response to the concerns posed by the close-sourced models, follow-up works suggest fine-tuning language models locally for evaluations. For instance, PandaLM Wang et al. (2024) constructs data based on Alpaca instructions and GPT-3.5 annotation, and then fine-tunes LLaMA-7B Touvron et al. (2023) as a judge model. JudgeLM Zhu et al. (2023b) constructs data from diversified instruction sets and GPT-4 annotations, and fine-tunes Vicuna Chiang et al. (2023) as a scalable judge model. Auto-J Li et al. (2023a) constructs judgment data upon multiple scenarios to train a generative judge model, which can provide both judgement and critic. Prometheus Kim et al. (2023) defines thousands of evaluation criteria, constructs a feedback dataset based on GPT-4, and finetunes a fine-grained judge model.
Despite relying on models with 7B or 13B parameters, these fine-tuned judge models all achieve comparable accuracy with GPT-3.5 or GPT-4 on their respective testsets. However, the evaluation is mostly conducted on the in-domain testsets constructed similarly with the training data. A thorough examination of the evaluation capability of the fine-tuned judges is in urgent demand.
@@ 3 How Far can Fine-tuned Judges Go?
In this section, we make a comprehensive empirical study based on four representative fine-tuned judge models as listed in Table 1.
@@ 3.1 Preliminary: Fine-tuned LLM Judge
Model JudgeLM-test PandaLM-test Auto-J-test Prometheus-test
accuracy F1 accuracy F1 agreement pearson-ind pearson-ood
JudgeLM-7B 82.39 72.97 68.17 65.18 45.3 0.398 0.384
PandaLM-7B 66.44 56.01 68.97 60.95 40.0 0.417 0.386
Auto-J-13B 77.79 62.64 72.17 64.10 53.6 0.614 0.591
Prometheus-13B 24.58 23.39 29.03 27.92 16.2 0.864 0.869
+grade-twice 54.24 50.04 45.25 43.58 47.8
GPT-3.5-0613 72.57 51.40 64.36 46.40 42.7 0.636 0.563
GPT-4-1106 84.24 72.83 75.78 71.51 56.9 0.742 0.743
Table 2: Results of evaluators on pairwise selection. Notice Prometheus can be transformed for pairwise selection by grading each answers and compare the scores, therefore we release both results with and without grade-twice.
3.2 Limitations of Fine-tuned Judges
3.2.1 Constrained by Evaluation Scheme
Model MT-Bench
@@ 3.2.1 Constrained by Evaluation Scheme
One of the most appealing attributes of LLMs is their generalization ability, enabling them to execute various tasks defined by various instructions Zhu et al. (2023a). Under the case of LLM evaluation, the instruction can also be formed in various schemes: pairwise selection, pointwise grading, etc. Since different judge models are fine-tuned on different schemes, we would like to verify their evaluation capability under the scheme defined by others. Specifically, we apply their publicly released checkpoints, and cross-validate the judge models on each other’s testsets [^2] We make minimal modifications to the predefined prompts to adapt the model to different schemes. For detailed prompts please refer to Appendix A.2..
As shown in Table 2, all four models perform the best on their own training schemes, respectively, with results comparable with GPT-4. However, if we employ a model on an evaluation scheme where it is not trained, the evaluation performance would drop by a large margin. For example, using a pairwise model for pointwise grading, or using a pointwise model for pairwise selection, would all lead to catastrophic performance degradation. On the contrary, close-sourced models such as GPT-3.5 or GPT-4 consistently exhibit superior performance across various evaluation schemes.
We also validate the judge models on MT-bench Zheng et al. (2023), which is a multi-turn meta-evaluation dataset\@footnotemark. As shown in Table 3, while the four models are all trained for single-turn evaluation, they underperform GPT-4 on MT-bench by a large margin. This demonstrates that the finetuned judge models are overfitted to their respective evaluation schemes and lost their generalizability.
3.2.2 Biased Towards Superficial Quality
Recently, there has been a lot of research on the bias of LLM-based evaluators, namely the evaluator would favor more verbose answers, or answers with similar format Wang et al. (2023b); Saito et al. (2023). Subsequently, Zeng et al. (2023) proposed LLMBar as a testbed for the fairness of evaluators. It comprises one natural testset (Natural) and four adversarial testsets (Neighbor, Manual, GPTOut, GPTInst), and the adversarial testsets consist of paired outputs with a correct answer and an incorrect answer with better superficial quality (e.g., more fluent, more verbose, etc.). We evaluate the judge models on LLMBar [^3] The detailed prompts are present in Appendix A.2. as shown in Table 4.
As can be seen, the fine-tuned judge models perform poorly on adversarial testsets, even worse than random-guess. This notifies that they are severely biased toward superficial quality such as formality or verbosity, while neglecting crucial properties such as instruction following, resulting in the preference for incorrect answers. On the other hand, GPT-4 does not over-rely on the superficial features and achieves decent accuracy on all the testsets. This notifies that the superior performance of fine-tuned judges on the in-domain testsets may rely on spurious features Niven and Kao (2019), instead of really differentiating response quality.
@@ 3.2.2 Biased Towards Superficial Quality
GPT-3.5-0613 57.50 57.10 62.60 60.27 72.10 72.08 95.10 80.80 95.54 61.70
GPT-4-1106 72.50 72.50 72.00 71.44 84.50 84.78 94.50 82.78 98.75 65.55
Table 5: Results of evaluators on aspect-specific evaluation. w/o adapt denotes using the original prompt without adaptation to the specific aspect. As HaluEval and ToxicChat are both binary classifications, we apply Auto-J and Prometheus with pointwise grading and conduct a grid search to determine the classification threshold. On the other hand, as SALAD-Bench is a pairwise classification, we apply pairwise selection models, namely JudgeLM, PandaLM, and Auto-J to select a better response.
We also fine-tune a DeBERTa-based judge model in a classification style on LLMBar (please refer to Section 3.3 for details). It deserves noticing that the DeBERTa-based evaluator also outperforms the LLM-based evaluator by a large margin in terms of fairness. This inspires us that the bias of LLM-based evaluators may come from the casual language modeling process. While the model is trained to generate fluent and verbose responses, it also tends to prefer fluent and verbose response when employed for evaluation.
3.2.3 Incapable of Aspect-specific Evaluation
LLM evaluation covers various aspects such as helpfulness, safety, etc., and sometimes we are particularly interested in a specific aspect. While previous work primarily assess the judge models from a general perspective, in this part, we would like to assess their evaluation capability on fine-grained aspects. We select the following three datasets:
@@ 3.2.3 Incapable of Aspect-specific Evaluation
    HaluEval Li et al. (2023b): This dataset is used for factuality evaluation. It consists of three domains: Question-Answering, Summary, and Dialogue.
@@
    ToxicChat Lin et al. (2023): This dataset is used for toxicity evaluation.
@@
    SALAD-Bench Li et al. (2024): This dataset is used for safety evaluation.
@@
3.2.4 No Benefits from Prompt Engineering
One of the most appealing features of LLM is it can benefit from delicate prompt engineering. Various strategies have been proposed to improve the LLM’s capability on various tasks, including text evaluation. In this section, we select two representative strategies and validate their effectiveness for fine-tuned judge models: